OpenStudy (anonymous):

For triangle ABC,angle A=35°angle B=55° and side c=70. find side a

OpenStudy (anonymous):

angle C is \(180-55-35=90\) so law of sines tell you \[\frac{a}{\sin(35)}=\frac{70}{\sin(90)}\]

OpenStudy (anonymous):

C=90 A/SIN a=C/SIN c

OpenStudy (anonymous):

since \(\sin(90)=1\) you get \[a=70\times \sin(35) \] then a calculator